1.)

Qb = 32 -

Consider the market for wheat as given in Question 1 above. Suppose government passes a regulation which makes it illegal for sellers to sell at a price higher than 30 cents per pound.

a.) What is total market surplus post-regulation? Compare it with pre-regulation market surplus.

b.) Calculate the deadweight loss created by the regulation.

c.) Calculate the value of consumer surplus post-regulation. Compare it with pre-regulation consumer surplus.

d.) Calculate the value of producer surplus post-regulation. Compare it with pre-regulation producer surplus.

e.) Identify other undesirable effects of this regulation.

--> If you can help me with this question that would be great.
Answer #1

Initially, demand = supply

or, 32-2/5P=4/5P-16

or, 6/5P=48

or, P=48*5/6 = $40

and Q=4/5P-16 = 4/5*40 - 16 = 16 units

From the demand and supply equations, demand curve has vertical intercept of 80 (when Q=0,P=80) and supply curve has vertical intercept of 20 (when Q=0,P=20)

Total surplus = consumer surplus + producer surplus = 1/2* (vertical intercept of demand curve-equilibrium price)*equilibrium quantity + 1/2*(equilibrium price - vertical intercept of supply curve)*equilibrium quantity = 1/2*(80-40)*16 + 1/2*(40-20)*16 = $320+$160 =$480

a) If government places restriction on price at $30, Quantity supplied = 4/5P-16 = 4/5*30 - 16 = 8 units.

Thus, new equilibrium quantity = 8 units

Now, when QD=8, P=$60

Then, total surplus = consumer surplus + producer surplus = 1/2*(sum of lengths of parallel sides)*height + 1/2*base*height = 1/2*{(80-30)+(60-30)}*8 + 1/2*(30-20)*8 = $320+$40 = $360

b) Dead weight loss = Initial total surplus - final total surplus = $480-$360 = $120

c) Consumer surplus remains the same i.e, $320

d) Producer surplus falls from $160 to $40 i.e, by $120.

e) Total surplus falls by $120, that means welfare of the society diminishes due to regulation.